# A Survey on Graph Neural Networks for Time Series: Forecasting, Classification, Imputation, and Anomaly Detection

### v1

@@ 3 Framework and Categorization
In this section, we present a comprehensive task-oriented taxonomy for GNNs within the context of time series analysis (subsection 3.1). Subsequently, we elucidate the foundational principles for encoding time series data across various tasks by introducing a unified methodological framework of GNN architectures (subsection 3.2). According to the framework, all architectures are composed of a similar graph-based processing module $f_{\theta}$ and a second module, $p_{\phi}$, specialized for the downstream task. Here, we also provide a general pipeline for analyzing time series data using GNNs. The combination of these perspectives offers a comprehensive overview of GNN4TS.
@@ 3.1 Task-Oriented Taxonomy
(d) Graph neural networks for time series classification: Convert green series classification into a graph (top) or node (bottom) classification task.
@@
In Figure 3, we illustrate a task-oriented taxonomy of GNNs encompassing the primary tasks and mainstream modeling perspectives for time series analysis, and showcasing the potential of GNN4TS. To summarize, our survey emphasizes four categories: Time series forecasting, anomaly detection, imputation, and classification. These tasks are performed on top of the time series representations learned by spatial-temporal graph neural networks (STGNNs), which serve as the foundation for encoding time series data in existing literature across various tasks. We detail this in subsection 3.2.
Time Series Forecasting. This task is centered around predicting future values of the time series based on historical observations, as depicted in 4a. Depending on application needs, we categorize this task into two types: single-step forecasting and multi-step forecasting. The former is meant to predict single future observations of the time series once at a time, i.e., the target at time $t$ is $\mathbf{Y}:=\mathbf{X}_{t+H}$ for $H\in\mathbb{N}$ steps ahead, while the latter makes predictions for a time interval, e.g., $\mathbf{Y}:=\mathbf{X}_{t+1:t+H}$. Solutions to both predictive cases can be cast in the optimization form:
@@
where $f_{\theta}(\cdot)$ and $p_{\phi}(\cdot)$ represent a spatial-temporal GNN and the predictor, respectively. Details regarding the $f_{\theta}(\cdot)$ architecture are given in subsection 3.2 while the predictor is, normally, a multi-layer perceptron. In the sequel, we denote by $\mathbf{X}_{t-T:t}$ and $\mathbf{A}_{t-T:t}$ a spatial-temporal graph $\mathcal{G}=\{\mathcal{G}_{t-T},\mathcal{G}_{t-T+1},\cdots,\mathcal{G}_{t}\}$ with length $T$. If the underlying graph structure is fixed, then $\mathbf{A}_{t}:=\mathbf{A}_{t-1}$. $\mathcal{L}_{\textsc{f}}(\cdot)$ denotes the forecasting loss, which is typically a squared or absolute loss function in most works, e.g., STGCN [66] and MTGNN [61]. Most existing works minimize the error between the forecasting and the ground truth $\mathbf{Y}$ through the Equation 6; this process is known as deterministic time series forecasting. Besides, we have probabilistic time series forecasting methods, such as DiffSTG [67], that share the same objective Equation 6 function though it is not directly optimized. Based on the size of the forecasting horizon $H$, we can conduct either short-term or long-term forecasting.
@@
Time Series Anomaly Detection. This task focuses on detecting irregularities and unexpected event in time series data (see 4b). Detecting anomalies requires determining when the anomalous event occurred, while diagnosing them requests gaining insights about how and why the anomaly occurred. Due to the general difficulty of acquiring anomaly events, current research commonly treats anomaly detection as an unsupervised problem that involves the design of a model describing normal, non-anomalous data. The learned model is then used to detect anomalies by generating a high score whenever an anomaly event occurs. The optimization process can be formulated as
$\theta^{*},\phi^{*}=\mathop{\arg\min}\limits_{{\theta,\phi}}\mathcal{L}_{\textsc{r}}\Big{(}p_{\phi}\big{(}f_{\theta}(\mathbf{X}_{t-T:t},\mathbf{A}_{t-T:t})\big{)},\mathbf{Y}\Big{)},$ (7)
where $f_{\theta}(\cdot)$ and $p_{\phi}(\cdot)$ denote the spatial-temporal GNN and predictor, respectively. Similar to the forecasting task, the detector can simply be a multi-layer perceptron. Commonly [48, 68], we optimize Equation 7 on non-anomalous training data to minimize the residual error between the input series and the predicted (reconstructed) one, where $\mathbf{Y}:=\mathbf{X}_{t}$; some methods intrinsically deal with training data possibly contaminated by anomalous observations [69]. Existing works [49, 70] also formulate this task as a one-step-ahead forecasting task by dropping the last observation as $\mathbf{X}_{t-T:t-1},\mathbf{A}_{t-T:t-1}$. In both cases, the reconstruction or forecast discrepancy should be large when provided with anomalous instances. The threshold discriminating between normal and anomalous data is a critical issue and should account for the fact that anomalies are typically rare events. It follows that the threshold is calibrated to align with a desired false alarm rate [71]. As it is important to diagnose the cause of anomaly events a common strategy envisages to design the statistics so that it computes discrepancies for each individual channel node before aggregating scores into a single anomaly score value [72]. In this way, the channel variables responsible for the anomaly events can be identified by computing their respective contributions to the final score.
@@
Time Series Imputation. This task is centered around estimating and filling in missing or incomplete data points within a time series dataset (4c). Current research in this domain can be broadly classified into two main approaches: In-sample imputation and out-of-sample imputation. In-sample imputation involves filling missing values in a given time series, while out-of-sample imputation pertains to inferring missing data not present in the training dataset. We formulate the learning objective as follows:
@@
Figure 5: Methodology-oriented taxonomy of graph neural network for time series analysis.
Time Series Classification. This task aims to assign a categorical label to a given time series based on its underlying patterns or characteristics. Rather than capturing patterns within a time series data sample, the essence of time series classification resides in discerning differentiating patterns that help separate samples based on their class labels. The objective of classifying a time series can be expressed as:
@@
where $f_{\theta}(\cdot)$ and $p_{\phi}(\cdot)$ denote a vanilla GNN and classifier to be learned, respectively. Using univariate time series classification as an example 4d, the task can be formulated as either a graph or node classification task. In the case of graph classification (Series2Graph) [73], each series is transformed into a graph, and the graph will input a GNN to generate a classification output. This can be achieved by dividing a complete series into multiple subsequences with a window size, $W$, serving as graph nodes, $\mathbf{X}\in\mathbb{R}^{N\times W}$, and an adjacency matrix, $\mathbf{A}$, describing the relationships between subsequences. A simple GNN, $f_{\theta}(\cdot)$, then employs graph convolution and pooling to obtain a condensed graph feature to be exploited by a classifier $p_{\phi}(\cdot)$ which assigns a class label to the graph. Alternatively, the node classification formulation (Series2Node), treats each series as a node in a dataset graph. Unlike Series2Graph, it constructs an adjacency matrix representing the relationships between multiple distinct series in a given dataset [74]. With several series of length $T$ stacked into a matrix $\mathbf{X}\in\mathbb{R}^{N\times T}$ as node features and $\mathbf{A}$ representing pairwise relationships, the GNN operation, $f_{\theta}(\cdot)$, aims at leveraging the relationships across different series for accurate node series classification [75]. In all cases, $\mathbf{Y}$ is typically a one-hot encoded vector representing the categorical label of a univariate or multivariate time series.
@@ 3.2 Unified Methodological Framework
  • Spatial Module. To model dependencies between time series over time, STGNNs employ the design principles of GNNs on static graphs. These can be further categorized into three types: Spectral GNNs, spatial GNNs, and a combination of both (i.e., hybrid) [29]. Spectral GNNs are based on spectral graph theory and use the graph shift operator (like the graph Laplacian) to capture node relationships in the graph frequency domain [76, 77, 28]. Differently, spatial GNNs simplify spectral GNNs by directly designing filters that are localized to each node’s neighborhood. Hybrid approaches combine both spectral and spatial methodologies to capitalize on the strengths of each method.
  • Temporal Module. To account for temporal dependencies in time series, STGNNs incorporate temporal modules that work in tandem with spatial modules to model intricate spatial-temporal patterns. Temporal dependencies can be represented in either the time or frequency domains. In the first category of methods, approaches encompass recurrence-based (e.g., RNNs [25]), convolution-based (e.g., TCNs [78]), attention-based (e.g., Transformers [27]), and a combination of these (i.e., hybrid). For the second category, analogous techniques are employed, followed by orthogonal space projections [28], such as the Fourier transform.
General Pipeline. In Figure 6, we showcase a general pipeline that demonstrates how STGNNs can be integrated into time series analysis. Given a time series dataset, we first process it using the data processing module, which performs essential data cleaning and normalization tasks, including the extraction of time series topology (i.e., graph structures). Subsequently, STGNNs are utilized to obtain time series representations, which can then be passed to different handlers (i.e., downstream task prediction module) to execute various analytical tasks, such as forecasting and anomaly detection.
@@ 4 GNNs for Time Series Forecasting
TABLE II: Summary of representative graph neural networks for time series forecasting. Task notation: The first letter, “M” or “S”, indicates multi-step or single-step forecasting, and the second letter, “S” or “L”, denotes short-term or long-term forecasting. Architecture notation: “D” and “C” represent “Discrete” and “Continuous”; “C” and “F” stand for “Coupled” and “Factorized”. Temporal module notation: “T” and “F” signify “Time Domain” and “Frequency Domain”; “R”, “C”, “A”, and “H” correspond to “Recurrence”, “Convolution”, “Attention”, and “Hybrid”. Input graph notation: “R” indicates that a pre-calculated graph structure (with a certain graph heuristic) is a required input of the model, “NR” that such graph is not required (not a model’s input), while “O” signifies that the model can optionally exploit given input graphs. Notation of learned graph relations: “S” and “D” indicate “Static” and “Dynamic”. Notation of adopted graph heuristics: “SP”, “PC”, “PS”, and “FD” denote “Spatial Proximity”, “Pairwise Connectivity”, “Pairwise Similarity”, and “Functional Dependency”, respectively. The “Missing Values” column indicates whether corresponding methods can handle missing values in input time series.
@@
Approach Year Venue Task Architecture Spatial Module Temporal Module Missing Values Input Graph Learned Relations Graph Heuristics DCRNN [80] 2018 ICLR M-S D-C Spatial GNN T-R No R - SP STGCN [66] 2018 IJCAI M-S D-F Spectral GNN T-C No R - SP ST-MetaNet [82] 2019 KDD M-S D-F Spatial GNN T-R No R - SP, PC NGAR [83] 2019 IJCNN S-S D-F Spatial GNN T-R No R - - ASTGCN [84] 2019 AAAI M-S D-F Spectral GNN T-H No R - SP, PC ST-MGCN [54] 2019 AAAI S-S D-F Spectral GNN T-R No R - SP, PC, PS Graph WaveNet [85] 2019 IJCAI M-S D-F Spatial GNN T-C No O S SP MRA-BGCN [86] 2020 AAAI M-S D-C Spatial GNN T-R No R - SP MTGNN [61] 2020 KDD S-S, M-S D-F Spatial GNN T-C No NR S - STGNN* [87] 2020 WWW M-S D-C Spatial GNN T-H No R - SP GMAN [88] 2020 AAAI M-S D-C Spatial GNN T-A No R - SP SLCNN [89] 2020 AAAI M-S D-F Hybrid T-C No NR S - STSGCN [90] 2020 AAAI M-S D-C Spatial GNN T No R - PC StemGNN [62] 2020 NeurIPS M-S D-F Spectral GNN F-C No NR S - AGCRN [91] 2020 NeurIPS M-S D-C Spatial GNN T-R No NR S - LSGCN [92] 2020 IJCAI M-S D-F Spectral GNN T-C No R - SP STAR [93] 2020 ECCV M-S D-F Spatial GNN T-A No R - PC GTS [64] 2021 ICLR M-S D-C Spatial GNN T-R No NR S - GEN [94] 2021 ICLR S-S D-F Spatial GNN T-R No R - - Z-GCNETs [95] 2021 ICML M-S D-C Spatial GNN T-C No NR S - STGODE [79] 2021 KDD M-S C-F Spatial GNN T-C No R - SP, PS STFGNN [57] 2021 AAAI M-S D-F Spatial GNN T-C No R - SP, PS DSTAGNN [96] 2022 ICML M-S D-F Spectral GNN T-H No R - PC, PS TPGNN [97] 2022 NeurIPS S-S, M-S D-F Spatial GNN T-A No NR D - MTGODE [22] 2022 IEEE TKDE S-S, M-S C-C Spatial GNN T-C No NR S - STG-NCDE [98] 2022 AAAI M-S C-C Spatial GNN T-C Yes NR S - STEP [99] 2022 KDD M-S D-F Spatial GNN T-A No NR S - Chauhan et al. [100] 2022 KDD M-S - - - Yes O S SP RGSL [101] 2022 IJCAI M-S D-C Spectral GNN T-R No R S SP, PC FOGS [102] 2022 IJCAI M-S - - - No NR S - METRO [103] 2022 VLDB M-S D-C Spatial GNN T No NR D - SGP [104] 2023 AAAI M-S D-F Spatial GNN T-R No R - SP Jin et al. [28] 2023 arXiv M-S, M-L D-F Spectral GNN F-H No NR S -
@@
Time series forecasting aims to predict future time series values based on historical observations. The origin of time series forecasting can be traced back to statistical auto-regressive models [105], which forecast future values in a time series based on a linear combination of its past values. In recent years, deep learning-based approaches have demonstrated considerable success in forecasting time series by capturing nonlinear temporal and spatial patterns more effectively [22]. Techniques such as recurrent neural networks (RNNs), convolutional neural networks (CNNs), and attention-based neural networks have been employed. However, many of these approaches, such as LSTNet [106] and TPA-LSTM [107], overlook and implicitly model the rich underlying dynamic spatial correlations between time series. Recently, graph neural network (GNN)-based methods have shown great potential in explicitly and effectively modeling spatial and temporal dependencies in multivariate time series data, leading to enhanced forecasting performance.
@@
GNN-based forecasting models can be categorized and examined from multiple perspectives. In terms of forecasting tasks, while many models focus on multi-step forecasting (i.e., predicting multiple consecutive steps ahead based on historical observations), a minority also discuss single-step forecasting (i.e., predicting the next or one arbitrary step ahead). From a methodological standpoint, these models can be dissected from three aspects: (1) Modeling spatial (i.e., inter-variable) dependencies, (2) modeling inter-temporal dependencies, and (3) the architectural fusion of spatial and temporal modules for time series forecasting. A summary of representative works is in Table II.

### v3

@@ 3 Framework and Categorization
In this section, we present a comprehensive task-oriented taxonomy for GNNs within the context of time series analysis (subsection 3.1). Subsequently, we investigate how to encode time series across various tasks by introducing a unified methodological framework for GNN architectures (subsection 3.2). According to the framework, all architectures are composed of a similar graph-based processing module $f_{\theta}$ and a second module $p_{\phi}$ specialized in downstream tasks.
@@ 3.1 Task-oriented Taxonomy
(d) Graph neural networks for time series classification: Formulates green series classification as a graph (top) or node (bottom) classification task.
@@
In Figure 3, we illustrate a task-oriented taxonomy of GNNs encompassing the primary tasks and mainstream modeling for time series analysis, and showcasing the potential of GNN4TS. This survey focuses on four categories: time series forecasting, anomaly detection, imputation, and classification. These tasks are performed on top of the time series representations learned by spatial-temporal graph neural networks (STGNNs), which serve as the foundation for encoding time series data in existing literature across various tasks. We detail this in subsection 3.2.
Time Series Forecasting. This task is centered around predicting future values of the time series based on historical observations, as depicted in 3a. Depending on application needs, we categorize this task into two types: single-step-ahead forecasting and multi-step-ahead forecasting. The former is meant to predict single future observations of the time series once at a time, i.e., the target at time $t$ is $\mathbf{Y}:=\mathbf{X}_{t+H}$ for some $H\in\mathbb{N}$ steps ahead, while the latter makes predictions for a time interval, e.g., $\mathbf{Y}:=\mathbf{X}_{t+1:t+H}$. Parameterized solutions to both predictive cases can be derived by optimizing
@@
where $f_{\theta}(\cdot)$ and $p_{\phi}(\cdot)$ represent a spatial-temporal GNN and the predictor, respectively. Details regarding the $f_{\theta}(\cdot)$ architecture are given in subsection 3.2 while the predictor is, normally, a multi-layer perceptron. In the sequel, we denote by $\mathbf{X}_{t-T:t}$ and $\mathbf{A}_{t-T:t}$ a spatial-temporal graph $\mathcal{G}=\{\mathcal{G}_{t-T},\mathcal{G}_{t-T+1},\cdots,\mathcal{G}_{t}\}$ with length $T$. If the underlying graph structure is fixed, then $\mathbf{A}_{t}:=\mathbf{A}$. $\mathcal{L}_{\textsc{f}}(\cdot)$ denotes the forecasting loss, which is typically a squared or absolute loss function, e.g., STGCN [54] and MTGNN [51]. Most existing works minimize the error between the forecasting and the ground truth $\mathbf{Y}$ through Equation 6; this process is known as deterministic time series forecasting. Besides, we have probabilistic time series forecasting methods, such as DiffSTG [55], that share the same objective Equation 6 function though it is not directly optimized. Based on the size of the forecasting horizon $H$, we end up in either short-term or long-term forecasting.
@@
Time Series Anomaly Detection. This task focuses on detecting irregularities and unexpected events in time series data (3b). Detecting anomalies requires determining when the anomalous event occurred, while diagnosing them requests gaining insights about how and why the anomaly occurred. Due to the general difficulty of acquiring anomaly events, current research commonly treats anomaly detection as an unsupervised problem that involves the design of a model describing normal, non-anomalous data. The learned model is then used to detect anomalies by generating a high score whenever an anomaly event occurs. This model learning process mirrors the forecasting optimization, Equation 6, with $f_{\theta}(\cdot)$ and $p_{\phi}(\cdot)$ denote the spatial-temporal GNN and the predictor, respectively. In general, the spatial-temporal GNN and the predictor are trained on normal, non-anomalous data using either forecasting [36, 56] or reconstruction [35, 57] optimization approaches, with the aim of minimizing the discrepancy between the normal input and the forecast (or reconstructed) series. Nonetheless, when these models are put to use for detecting anomalies, they are expected to fail in minimizing this discrepancy upon receiving anomalous input. This inability to conform to the expected low-discrepancy model behavior during anomaly periods creates a detectable difference, facilitating the detection of anomalies. The threshold separating normal and anomalous data is a sensitive hyperparameter that should be set considering the rarity of anomalies and aligned with a desired false alarm rate [58]. Lastly, to diagnose the causes of anomalies, a common strategy involves calculating discrepancies for each channel node and consolidating these into a single anomaly score [59]. This approach allows for the identification of the channel variables responsible for the anomaly events by calculating their respective contributions to the final score.
@@
Time Series Imputation. This task is centered around estimating and filling in missing or incomplete data points within a time series (3c). Current research in this domain can be broadly classified into two main approaches: in-sample imputation and out-of-sample imputation. In-sample imputation involves filling missing values in a given time series, while out-of-sample imputation pertains to inferring missing data not present in the training dataset. We formulate the learning objective as follows:
@@
{forest}
forked edges, for tree= fill=level0!80, grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=8.2em,font=,fill=level1!80,, where level=2text width=6.5em,font=,fill=level2!80,, where level=3text width=9.5em,font=,fill=level3!60,, [GNNs for Time Series Analysis [Overall Architecture [Discrete [Factorized [Forecasting: [54],[41],[50],[44],[60],[61], leaf, text width=18.8em] [Others: [62],[63],[64],[65],[66],[67],[38], leaf, text width=19em] ] [Coupled [Forecasting: [68],[69],[70],[52],[71],[72], leaf, text width=18.8em] [Others: [36],[56],[33],[73],[37], leaf, text width=15em] ] ] [Continuous [Factorized [Forecasting and other tasks: [74],[57], leaf, text width=17em] ] [Coupled [Forecasting: [23],[75], leaf, text width=10em] ] ] ] [Temporal Module [Time Domain [Recurrence-based [Forecasting: [68],[76],[41],[69],[49],[77], leaf, text width=18.7em] [Others: [78],[79],[56],[80],[37],[81], leaf, text width=17em] ] [Convolution-based [Forecasting: [54],[48],[51],[71],[44],[23], leaf, text width=19em] [Others: [82],[62],[63],[64],[65],[66], leaf, text width=17em] ] [Attention-based [Forecasting: [70],[83],[60],[61], leaf, text width=15em] [Others: [35],[84],[33],[67],[85],[38], leaf, text width=16.5em] ] [Hybrid [Forecasting: [86],[87],[88], leaf, text width=12em] [Anomaly Detection: [89],[90], leaf, text width=14em] ] ] [Frequency Domain [Convolution-based [Forecasting: [50],[91], leaf, text width=10em] ] [Hybrid [Forecasting: [28], leaf, text width=8em] ] ] ] [Spatial Module [Spatial GNNs [Message passing-based [Forecasting: [76],[92],[49],[74],[44],[93], leaf, text width=19em] [Others: [35],[36],[79],[33],[64],[65], leaf, text width=17em] ] [Graph diffusion-based [Forecasting: [68],[48],[52],[94], leaf, text width=14.5em] [Imputation: [73], leaf, text width=8em] ] ] [Spectral GNNs [Polynomial approximation-based [Forecasting: [54],[86],[50],[88],[72],[28], leaf, text width=18.6em] ] ] [Hybrid [Forecasting: [95], leaf, text width=8em] ] ] ]
Figure 5: Methodology-oriented taxonomy of graph neural networks for time series analysis.
Time Series Classification. This task aims to assign a categorical label to a given time series based on its underlying patterns or characteristics. Rather than capturing patterns within a time series data sample, the essence of time series classification resides in discerning differentiating patterns that help separate samples based on their class labels. The optimization problem can be expressed as:
@@
where $f_{\theta}(\cdot)$ and $p_{\phi}(\cdot)$ denote, e.g., a GNN and a classifier to be learned, respectively. Using univariate time series classification as an example 3d, the task can be formulated as either a graph or node classification task. In the case of graph classification (Series-as-Graph) [96], each series is transformed into a graph, and the graph will be the input of a GNN to generate a classification output. This can be achieved by dividing a series into multiple subsequences with a window size, $W$, serving as graph nodes, $\mathbf{X}\in\mathbb{R}^{N\times W}$, and an adjacency matrix, $\mathbf{A}$, describing the relationships between subsequences. A simple GNN, $f_{\theta}(\cdot)$, then employs graph convolution and pooling to obtain a condensed graph feature to be exploited by a classifier $p_{\phi}(\cdot)$ which assigns a class label to the graph. Alternatively, the node classification formulation (Series-as-Node), treats each series as a node in a dataset graph. Series-as-Node constructs an adjacency matrix representing the relationships between multiple distinct series in a given dataset [63]. With several series of length $T$ stacked into a matrix $\mathbf{X}\in\mathbb{R}^{N\times T}$ as node features and $\mathbf{A}$ representing pairwise relationships, the GNN operation, $f_{\theta}(\cdot)$, aims at leveraging the relationships across different series for accurate node series classification [97]. In all cases, $\mathbf{Y}$ is typically a one-hot encoded vector representing the categorical label of a univariate or multivariate time series.
@@ 3.2 Unified Methodological Framework
Spatial Module. To model dependencies between time series over time, STGNNs employ the design principles of GNNs on static graphs. These can be further categorized into three types: spectral GNNs, spatial GNNs, and a combination of both (i.e., hybrid) [29]. Spectral GNNs are based on spectral graph theory and use the graph shift operator (like the graph Laplacian) to capture node relationships in the graph frequency domain [98, 99, 28]. Differently, spatial GNNs simplify spectral GNNs by directly designing filters that are localized to each node’s neighborhood. The approaches in this category can be broadly classified into two types: graph diffusion-based and message passing-based. Notably, graph transformers [100] represent a specialized extension of message passing neural networks [101], further expanding the capabilities of this paradigm. Hybrid approaches combine both spectral and spatial methodologies to capitalize on the strengths of each method.
Temporal Module. To account for temporal dependencies in time series, STGNNs incorporate temporal modules that work in tandem with spatial modules to model intricate spatial-temporal patterns. Temporal dependencies can be represented in either the time or frequency domains. The approaches in the first category encompass recurrence-based (e.g., RNNs [26]), convolution-based (e.g., TCNs [102]), attention-based (e.g., transformers [27]), and a combination of these (i.e., hybrid). For the second category, analogous techniques are employed, followed by orthogonal space projections [28], such as the Fourier transform.
@@
General Pipeline. In Figure 6, we showcase a general pipeline that shows how STGNNs can be integrated into time series analysis. Given a time series dataset, we first process it using the data processing module, which performs essential data cleaning and normalization tasks, including the extraction of time series topology (i.e., graph structures). Subsequently, STGNNs are utilized to obtain time series representations, which can then be passed to different handlers (i.e., downstream task prediction module) to execute various analytical tasks, such as forecasting and anomaly detection.
@@ 4 GNNs for Time Series Forecasting
TABLE II: Summary of representative graph neural networks for time series forecasting. Task notation: The first letter, “M” or “S”, indicates multi-step or single-step forecasting, and the second letter, “S” or “L”, denotes short-term or long-term forecasting. Architecture notation: “D” and “C” represent “Discrete” and “Continuous”; “C” and “F” stand for “Coupled” and “Factorized”. Temporal module notation: “T” and “F” signify “Time” and “Frequency” domains; “R”, “C”, “A”, and “H” correspond to “Recurrence”, “Convolution”, “Attention”, and “Hybrid”. Input graph notation: “R” indicates that a pre-calculated graph structure (with a certain graph heuristic) is a required input of the model, “NR” that such graph is not required (not a model’s input), while “O” signifies that the model can optionally exploit given input graphs. Notation of learned graph relations: “S” and “D” indicate “Static” and “Dynamic”. Notation of adopted graph heuristics: “SP”, “PC”, “PS”, and “FD” denote “Spatial Proximity”, “Pairwise Connectivity”, “Pairwise Similarity”, and “Functional Dependency”. The “Missing Values” column indicates whether corresponding methods can handle missing values in input time series.
@@
Approach Year Venue Task Architecture Spatial Module Temporal Module Missing Values Input Graph Learned Relations Graph Heuristics DCRNN [68] 2018 ICLR M-S D-C Spatial GNN T-R No R - SP STGCN [54] 2018 IJCAI M-S D-F Spectral GNN T-C No R - SP ST-MetaNet [76] 2019 KDD M-S D-F Spatial GNN T-R No R - SP, PC NGAR [104] 2019 IEEE IJCNN S-S D-F Spatial GNN T-R No R - - ASTGCN [86] 2019 AAAI M-S D-F Spectral GNN T-H No R - SP, PC ST-MGCN [41] 2019 AAAI S-S D-F Spectral GNN T-R No R - SP, PC, PS Graph WaveNet [48] 2019 IJCAI M-S D-F Spatial GNN T-C No O S SP MRA-BGCN [69] 2020 AAAI M-S D-C Spatial GNN T-R No R - SP MTGNN [51] 2020 KDD S-S, M-S D-F Spatial GNN T-C No NR S - STGNN* [87] 2020 WWW M-S D-C Spatial GNN T-H No R - SP GMAN [70] 2020 AAAI M-S D-C Spatial GNN T-A No R - SP SLCNN [95] 2020 AAAI M-S D-F Hybrid T-C No NR S - STSGCN [92] 2020 AAAI M-S D-C Spatial GNN T No R - PC StemGNN [50] 2020 NeurIPS M-S D-F Spectral GNN F-C No NR S - AGCRN [49] 2020 NeurIPS M-S D-C Spatial GNN T-R No NR S - LSGCN [105] 2020 IJCAI M-S D-F Spectral GNN T-C No R - SP STAR [83] 2020 ECCV M-S D-F Spatial GNN T-A No R - PC GTS [52] 2021 ICLR M-S D-C Spatial GNN T-R No NR S - GEN [106] 2021 ICLR S-S D-F Spatial GNN T-R No R - - Z-GCNETs [71] 2021 ICML M-S D-C Spatial GNN T-C No NR S - STGODE [74] 2021 KDD M-S C-F Spatial GNN T-C No R - SP, PS STFGNN [44] 2021 AAAI M-S D-F Spatial GNN T-C No R - SP, PS DSTAGNN [88] 2022 ICML M-S D-F Spectral GNN T-H No R - PC, PS TPGNN [60] 2022 NeurIPS S-S, M-S D-F Spatial GNN T-A No NR D - MTGODE [23] 2022 IEEE TKDE S-S, M-S C-C Spatial GNN T-C No NR S - STG-NCDE [75] 2022 AAAI M-S C-C Spatial GNN T-C Yes NR S - STEP [61] 2022 KDD M-S D-F Spatial GNN T-A No NR S - Chauhan et al. [107] 2022 KDD M-S - - - Yes O S SP RGSL [72] 2022 IJCAI M-S D-C Spectral GNN T-R No R S SP, PC FOGS [108] 2022 IJCAI M-S - - - No NR S - METRO [93] 2022 VLDB M-S D-C Spatial GNN T No NR D - SGP [77] 2023 AAAI M-S D-F Spatial GNN T-R No R - SP, PS HiGP [109] 2023 arXiv M-S D-F Spatial GNN T-R No R S SP, PS Jin et al. [28] 2023 arXiv M-S, M-L D-F Spectral GNN F-H No NR S - CaST [91] 2023 NeurIPS M-S D-F Spectral GNN T&F-C No O S PC GPT-ST [110] 2023 NeurIPS M-S D-F Spatial GNN T-C No NR S -
@@
Time series forecasting aims to predict future time series values based on historical observations. While deep learning models have demonstrated considerable success in forecasting time series by capturing nonlinear temporal and spatial patterns more effectively than the linear counterpart [23], many of these approaches, such as LSTNet [111] and TPA-LSTM [112], overlook and implicitly model the rich underlying dynamic spatial correlations between time series. Recently, graph neural network (GNN)-based methods have shown great potential in explicitly and effectively modeling spatial and temporal dependencies in multivariate time series data, leading to enhanced forecasting performance.
@@
GNN-based forecasting models can be categorized and examined from multiple perspectives. In terms of forecasting tasks, while many models focus on multi-step forecasting (i.e., predicting multiple consecutive steps ahead based on historical observations), a minority also discuss single-step forecasting (i.e., predicting the next or one arbitrary step ahead). From a methodological standpoint, these models can be dissected from three aspects: (1) modeling spatial (i.e., inter-variable) dependencies, (2) modeling inter-temporal dependencies, and (3) the fusion of spatial and temporal modules for time series forecasting. A summary of representative works is given in Table II.